# On the Robustness of Verbal Confidence of LLMs in Adversarial Attacks

### v1

@@ Appendix I Confidence Expressed Via Likert Scale
In this section, we plot the distribution of confidence scores for different models and CEMs and attempt to see whether it matches distributions of percentages in real world data. We plot the full confidence scores obtained from running all of our attacks against user queries for GPT-3.5 in Figures 4, 5, 6, 7, and for Llama-3-8B in Figures 8, 9, 10, 11. In general, we observe highly overconfident distributions, particularly for the Base methods and when using GPT-3.5 as a whole, where lower confidence scores below 50% are rarely predicted by the model, and even under adversarial attacks the confidence scores rarely get shifted to the lower levels of confidence.
To model a real-world data distribution, we choose the popular RACE dataset [31], a large scale multiple choice dataset that closely mirrors the use-case for our main experiments, and plot the distribution of mentions of percentages (e.g., 50%) in the dataset in Figure 12. We see that the distribution in real-world datasets does not match the model’s prediction. As such, the behaviour of the models is not a direct result of the natural distribution of mentions of percentages in related data, which we particularly see with the infrequence if 95% in real data compared to being the dominant model prediction. Hence, the overconfident and limited distribution (and as a result the invariance under attacks) is inherent in the model behaviour but is due to other factors. We test an additional dataset, Reddit Corpus (small) [22], in the conversational domain in Figure 13 to observe whether there are significant differences. When plotting percentages we again observe a wider and more uneven distribution compared to the LLMs. We lastly use the WikiText dataset [40] to model a generic distribution that is commonly used to train LLMs, and the results can be seen in Figure 14. The distribution is very similar to RACE.
In contrast to simply finding the distribution of percentages, we also examine how common phrases like Xsure or Yconfident are in the data. Using the Reddit Corpus (small) since it contains natural conversations, we plot the distribution of these phrases in Figure 15. Unlike the previous examples, we note that this distribution is more overconfident and resembles the skewed distributions of the LLMs. This signals that models are likely copying natural human conversational patterns of exaggeration rather than a balanced distribution of confidence scores as one would find and expect in real-world data.
@@
Figure 7: Confidence score distribution using GPT-3.5 with the SC method.
@@
Figure 11: Confidence score distribution using Llama-3-8B with the SC method.
@@
Calibration is a key aspect of safety-critical systems, requiring that model confidence predictions are closely aligned with the level of predictive performance. Calibration error, which results from a mismatch between the accuracy of the model on sets of data at given levels of confidence, is quantified using a degree of metrics. For this analysis, we use the widely used binning-based metric, expected calibration error (ECE) [46], to measure the degree of calibration error induced post-attack. We centre this analysis on the results for GPT-4 and Llama-3 70B given their very high level of performance and hence larger capability of being miscalibrated through adversarial attacks leading to lowered confidence scores.
Table 16 compares how the ECE is affected post TB and SSR attacks across the main four tested CEMs. A value of 10 was used for the number of bins. We can see how the models have very high calibration both before and after undergoing attacks, and that in most cases for GPT4 there is a increase in ECE on adversarial examples. In contrast, ECE is more mixed with Llama-3-70B with inconsistent patterns of miscalibration across different cases. This likely means that Llama is more overconfident and hence decreasing confidence through attacks can appear to reduce ECE. Nevertheless, in most cases the impact is significant between the original and adversarial scenarios, and thus warrants close attention. Figures 16 and 17 both show calibration diagrams for GPT4 and Llama-3-70B respectively to reveal the general miscalibration patterns across all CEMs and attack types. Overall, as expected, the adversarial data has a high mismatch between accuracy and confidence in low confidence bins as a whole.
@@
Although the miscalibration is generally high prior to the confidence attacks, we stress that evaluating the robustness of these models’ confidence behaviour is crucial. Despite most CEMs producing verbal confidence that do not have ideal calibration, previous work has shown that LLM generated confidence scores do follow relative ordering by demonstrating that there is a statistically significant difference in LLMs (such as the tested GPT-4 and Llama-3-70B models here) verbal confidence level on correct answers versus incorrect answers [44] whereby correct answers contain a higher level of confidence on average. By attacking confidence, we can affect this ordering and decouple LLMs relative preference in confidence to true answers, which highlights issues in model honesty and perception of veracity. In sum, if we wish to create well-calibrated models through methods such as recalibration, we must also consider how vulnerable they are to adversarial attacks and address these vulnerabilities as well otherwise we cannot ensure that models will be calibrated across all realistic inputs.
In addition, real-world production models and methods we detailed previously that currently using verbal confidence in industry are vulnerable irrespective of calibration and thus it is meaningful to study the impact on the verbal confidence. Furthermore, we emphasize that we also test CEMs designed to improve calibration, most notably MS, and our results show that while MS is far less overconfident and better calibrated it is more vulnerable to attacks in terms of error rate and changes in confidence, hence better calibrated models are no less vulnerable than their miscalibrated versions which we wish to portray. Lastly, we note that many important works such as in Guo et al. [21] have shown that a large amount of vision models and LLMs are heavily miscalibrated and despite this many have studied adversarial attacks against these models, both traditional and confidence-based such as in Obadinma et al. [43].
@@
Appendix O Confidence Score Change when Model Predicts Same Answer
@@
Ty. 4.4 3.5 6.2 3.1
[…]
VCA-TB 0.8 1.1 9.4 6.4
SSR 1.2 3.9 12.1 6.0
Ty. 1.6 9.2 4.8 11.1
We analyze how confidence scores change post attack on queries in cases when the model predicts the same answer compared to the original (an unsuccessful conventional adversarial attack) to ascertain how much damage occurs in these cases, focusing on whether reduction to confidence scores is still great in cases where the overall adversarial attack is unsuccessful. In Table 18 we run this analysis over all of the main datasets and attacks and CEMs for our Llama-3-8B and GPT-3.5, focusing on attacks on user questions. The results reveal that there is large variance between the Δ Aff. Cf for different methods, and that the difference can be stark (above 50%) to almost negligible. In most cases, the Δ Aff. Cf for these examples where the answers were unchanged is smaller than the overall Δ Aff. Cf when considering all samples seen in Table 4 such as in the case of TruthfulQA with Llama 3 using the Base method (0.2 vs. 19.3), meaning that the examples where the label is flipped are more susceptible to significant confidence reduction.
Appendix P Attacks Optimized for Pure Confidence Degradation
@@
As jailbreak-based attacks cannot guarantee that original answers are preserved on downstream examples, we choose the most effective perturbation-based attack method, SSR and change the attack objectives so that effective perturbations cannot also change the original generated answer, and observe how the difference in average confidence change is affected. Table 19 shows the results of these experiments. We observe how in some cases it appears allowing the attack algorithm to change the answer leads to larger drops in confidence, while in other cases the average confidence level on adversarial examples is higher than before. SC in particular seems to fare poorer when not allowed to induce label flips with the majority of cases the confidence drop is lower than the original attack algorithm. In all cases though, the difference is minor between both variants of the attack algorithm, showing that confidence vulnerability is not strongly tied to whether the model’s answer is changed as a result of adversarial perturbations.
Table 19: Attack performance using SSR when attacks do not lead to label flips.
@@
Appendix Q Detailed Defence Results Against ConfidenceTriggers
Table 20: Perplexity filter and LLM-Guard results against ConfidenceTriggers-AutoDAN on Llama-3-8B. A negative Δ shows average confidence increased filter.
@@
We test methods from the three main categories of jailbreak-based methods: Self-Processing Defences, Additional Helper Defences, and Input Permutation Defences [68]. Firstly, we test the popular the self-processing perplexity filter defence approach from Jain et al. [26]. Here, we optimize a perplexity threshold for each dataset/CEM combination that results in no normal samples being filtered but can filter any text with abnormal perplexity (such as those containing trigger prompts). We also test an additional helper defence using the LLM-Guard toolkit [19], using a detection based approach to filter out prompts with a high level of gibberish text. Finally, we include two input permutation defences, the Paraphrase defence [26], and a modified version of SmoothLLM [55]. The former uses an additional LLM (GPT-3.5) to rephrase an input prompt containing a trigger phrase. The latter samples multiple variants of an input prompt and performs permutations (character swapping) before feeding the modified prompts to the LLM and using majority voting to decide on the answer and averaging the confidence across the different prompts to attain the final confidence.
@@
We conduct our experiments on the same set of data as Table 4. We test mitigation strategies using deterministic CEMs (Base, CoT, and MS) since they are single sample and the minimal randomness allows us to determine clearly what strategies are most effective. Table 20 presents the results of the perplexity filter defence. We show the percentage of user queries that get filtered (%filtered) (only including those that the trigger originally led to a confidence decrease) and the resulting loss in average confidence on those samples (Δ conf) and %MI which refers to the percentage of benign prompts not containing the trigger phrase that would get filtered out. We can see that in most cases the majority of these samples do not get filtered and the samples that do not lead to a big drop in average confidence (so mainly queries with low confidence drops get filtered). Similarly, Table 20 presents the results for LLM-Guard. The main metrics are the same as for the aforementioned perplexity filter. Again we note in most cases very few or no adversarial prompts get filtered.
@@
Table 6 presents the results for the two input permutation defences. We show the difference in average confidence provided by the LLM between the original set of data and those with the permuted prompts without any trigger phrase (Δ conf O.) and with it (Δ conf A.). Additionally, we note what percentage of permuted examples with originally benign (Aff.) and trigger-included prompts (Adv Aff.) had an altered generated confidence level compared to the original un-permuted version. Finally, we note the average decrease in confidence after input permutation on the trigger-included prompts (Δ Adv). Our primary findings are that the permutation-based defences lead to significant changes in confidence both when the original prompt does and does not include a trigger phrase (with differences of up to  43%). The majority of examples in most cases have their confidence altered as a result of these defences and there are very few cases were minimal changes of confidence occur after permutation of an adversarial prompt signalling that these defences are largely ineffective.
@@
Nevertheless, without even considering defenses such as perplexity-based ones that are popular for trigger-based attacks, ConfidenceTriggers can still present issues in specific applications. For example, if an adversary gains access to a custom LLM and supply a compromised system prompt (or demonstration) with ConfidenceTriggers-AutoDAN that is difficult to notice, then an external user using this model with their own queries would be affected. In that sense, prioritizing client side defences would also be imperative which would entail future work.
@@
Δ conf O. Δ conf A. % Aff. % Adv Aff. Δ Adv. Δ conf O. Δ conf A. % Aff. % Adv Aff. Δ Adv. Δ conf O. Δ conf A. % Aff. %Adv Aff. Δ Adv.
@@
% filtered Δ conf %MI % filtered Δ conf %MI % filtered Δ conf %MI
@@
% filtered Δ conf % filtered Δ conf % filtered Δ conf
@@
Appendix R Jailbreak Defence Descriptions
Perplexity Filter [26]
We use a naive filter based on the text perplexity of the system prompt. The text perplexity of sequence of tokens is defined as the averaged negative log likelihood of the tokens. The log perplexity ($ppl$) is thus defined for a whole sequence of tokens $\mathbf{\mathcal{X}}=\{x_{1},\dots,x_{N}\}$ by $log(ppl)=\frac{-1}{|\mathbf{X}|}\sum_{i}P_{i}\;\mbox{log p}(x_{i}|x_{0:i-1})$. Given that the log likelihood of a token in a sequence represents its probability of appearing in that context, a high negative log likelihood would mean that the tokens in a sequence are likely unnatural or manipulated. For our implementation, we set a threshold such that any full prompt with a perplexity higher than the threshold is considered filtered. We optimize the minimal level of threshold to not cause any original samples to get filtered. Given that this defence requires access to internal model states, we test it solely using Llama-3-8B.
LLM-Guard [19]
@@
We follow the approach of Jain et al. [26] which follows the protocol of Kirchenbauer et al. [29] for this defence. We instruct GPT-3.5 to paraphrase the system prompt according to the instruction “Please paraphrase the following sentences in under 200 words and only output the revised version, the sentence is:”. We use a temperature of 0.7 for the paraphrased output and limit the number of tokens to 400. The goal is that by paraphrasing the impact of a system prompt containing a trigger is minimized.
SmoothLLM-Confidence-Attack [55]
@@
Given that the original method is formulated for jailbreaks, we make a few modifications for the defence to be better suited against ConfidenceTriggers. The modified SmoothLLM algorithm for defending against ConfidenceTriggers can be found in Algorithm 6. Essentially, we make $n$ copies of the system prompts, and perform random character perturbations on each (to $q\%$ of the characters), and then the outputs (generated answer and confidence) for each are aggregated. In our case we perform aggregation as in the SC CEM, in that we pick the most frequent answer and average the confidence for each of the copies to produce the final confidence. Given that there is no way to know if the confidence has been affected, averaging is the most straightforward aggregation strategy to employ. For the RandomPermutation() function in the algorithm, we utilize random character swaps in the prompt as it was found to be the most effective. A $q$ of indices in the string are chosen and are each replaced with a random character uniformly sampled from a 100-character alphabet derived from the python command string.printable. Note we do not use the threshold $\gamma$ in our version of the algorithm. We set $\delta=5$ and $q=0.1$ for our experiments.
@@
6:    $\mathcal{C}_{j}=\mbox{Get Confidence Score From LLM Using CEM With Input }\mathbf{Q}_{j}$
@@
9:$\mathcal{A}\leftarrow$ Mode($[\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{n}]$)
Appendix S ConfidenceTriggers Transferability
We test the transferability of triggers obtained using ConfidenceTriggers to determine whether their effectiveness is able to be applied to different datasets after being optimized for a single domain. We focus on the transfer of ConfidenceTriggers-AutoDAN triggers since they are more generalizable and in natural language and hence have higher potential to transfer across different datasets. We use the triggers optimized on TruthfulQA using GPT-3.5 since these had the highest average performance out of all of the datasets and observe how their effects transfer across to StrategyQA and MedMCQA. Table 24 shows the results of the transfer attacks. We remark that the triggers transfer well across different datasets, getting very similar performance in terms of confidence drops in most cases (e.g., 9.0 Δ Overall Cf. for the original versus 8.8 and 9.9 for the transfers for the Base CEM). In this sense, we can see that an optimized set of triggers can be effective across many different domains, heighting their danger.
@@
Δ Overall Cf. % Aff. Δ Aff. Cf %Flip(OC) %Flips(OW)
SQA Base 8.8 72.0 12.2 10.1 18.8
COT 2.2 39.5 5.6 4.8 18.9
MS 11.0 77.0 14.3 5.8 23.8
SC 15.0 79.5 18.9 33.8 36.7
Avg. 9.3 67.0 12.7 13.6 24.5

### v3

@@ Appendix I Confidence Expressed Via Likert Scale
In this section, we plot the distribution of confidence scores for different models and CEMs and attempt to see whether it matches distributions of percentages in real world data. We plot the full confidence scores obtained from running all of our attacks against user queries for GPT-3.5 in Figures 5, 6, 7, 8, and for Llama-3-8B in Figures 9, 10, 11, 12. In general, we observe highly overconfident distributions, particularly for the Base methods and when using GPT-3.5 as a whole, where lower confidence scores below 50% are rarely predicted by the model, and even under adversarial attacks the confidence scores rarely get shifted to the lower levels of confidence.
To model a real-world data distribution, we choose the popular RACE dataset [lai-etal-2017-race], a large scale multiple choice dataset that closely mirrors the use-case for our main experiments, and plot the distribution of mentions of percentages (e.g., 50%) in the dataset in Figure 13. We see that the distribution in real-world datasets does not match the model’s prediction. As such, the behaviour of the models is not a direct result of the natural distribution of mentions of percentages in related data, which we particularly see with the infrequence if 95% in real data compared to being the dominant model prediction. Hence, the overconfident and limited distribution (and as a result the invariance under attacks) is inherent in the model behaviour but is due to other factors. We test an additional dataset, Reddit Corpus (small) [henderson-etal-2019-repository], in the conversational domain in Figure 14 to observe whether there are significant differences. When plotting percentages we again observe a wider and more uneven distribution compared to the LLMs. We lastly use the WikiText dataset [merity2016pointer] to model a generic distribution that is commonly used to train LLMs, and the results can be seen in Figure 15. The distribution is very similar to RACE.
In contrast to simply finding the distribution of percentages, we also examine how common phrases like X% sure or Y% confident are in the data. Using the Reddit Corpus (small) since it contains natural conversations, we plot the distribution of these phrases in Figure 16. Unlike the previous examples, we note that this distribution is more overconfident and resembles the skewed distributions of the LLMs. This signals that models are likely copying natural human conversational patterns of exaggeration rather than a balanced distribution of confidence scores as one would find and expect in real-world data.
@@
Figure 8: Confidence score distribution using GPT-3.5 with the SC-CA method.
@@
Figure 12: Confidence score distribution using Llama-3-8B with the SC-CA method.
@@
Calibration is a key aspect of safety-critical systems, requiring that model confidence predictions are closely aligned with the level of predictive performance. Calibration error, which results from a mismatch between the accuracy of the model on sets of data at given levels of confidence, is quantified using a degree of metrics. For this analysis, we use the widely used binning-based metric, expected calibration error (ECE) [pakdaman2015ece], to measure the degree of calibration error induced post-attack. We centre this analysis on the results for GPT-4o and Llama-3-70B given their very high level of performance and hence larger capability of being miscalibrated through adversarial attacks leading to lowered confidence scores.
Table 18 compares how the ECE is affected post TB and SSR attacks across the main four tested CEMs. A value of 10 was used for the number of bins. We can see how the models have very high calibration both before (O.) and after (Adv.) undergoing attacks, and that in most cases for GPT-4o there is a increase in ECE on adversarial examples. In contrast, ECE is more mixed with Llama-3-70B with inconsistent patterns of miscalibration across different cases. This likely means that Llama is more overconfident and hence decreasing confidence through attacks can appear to reduce ECE. Nevertheless, in most cases the impact is significant between the original and adversarial scenarios, and thus warrants close attention. Figures 17 and 18 both show calibration diagrams for GPT-4o and Llama-3-70B respectively to reveal the general miscalibration patterns across all CEMs and attack types. Overall, as expected, the adversarial data has a high mismatch between accuracy and confidence in low confidence bins as a whole.
@@
Although the miscalibration is generally high prior to the confidence attacks, we stress that evaluating the robustness of these models’ confidence behaviour is crucial. Despite most CEMs producing verbal confidence that do not have ideal calibration, previous work has shown that LLM generated confidence scores do follow relative ordering by demonstrating that there is a statistically significant difference in LLMs (such as the tested GPT-4o and Llama-3-70B models here) verbal confidence level on correct answers versus incorrect answers [Omar2024.08.11.24311810] whereby correct answers contain a higher level of confidence on average. By attacking confidence, we can affect this ordering and decouple LLMs relative preference in confidence to true answers, which highlights issues in model honesty and perception of veracity. In sum, if we wish to create well-calibrated models through methods such as recalibration, we must also consider how vulnerable they are to adversarial attacks and address these vulnerabilities as well, otherwise we cannot ensure that models will be calibrated across all realistic inputs.
In addition, real-world production models and methods we detailed previously that are currently using verbal confidence in industry are vulnerable irrespective of the calibration. Furthermore, we emphasize that we also test CEMs designed to improve calibration, most notably MS, and our results show that while MS is far less overconfident and better calibrated in most scenarios, it is more vulnerable to attacks in terms of error rate and changes in confidence, hence better calibrated models are no less vulnerable than their miscalibrated versions. Lastly, we note that many important works such as in 10.5555/3305381.3305518 have shown that a large amount of vision models and LLMs are heavily miscalibrated and despite this many have studied adversarial attacks against these models, both traditional and confidence-based such as in obadinma2024calibration.
@@
Appendix O Brier Score Analysis
Table 20: Brier Score Results using the different types of attacks (against user queries) on SQA. Note in the majority of cases ( 80%) the confidence attacks increase the Brier score, showing miscalibration is induced. In most cases, the Brier score pre-attack is not severe (e.g., around 0.2) showing that verbal confidence is not heavily miscalibrated originally.
VCA-TF VCA-TB SSR Ty.
Pre Post Pre Post Pre Post Pre Post
Llama-3-8B SQA Base 0.51 0.52 0.51 0.51 0.51 0.42 0.51 0.42
CoT 0.07 0.15 0.07 0.19 0.07 0.20 0.07 0.28
MS 0.26 0.30 0.26 0.29 0.26 0.32 0.26 0.31
SC-CA 0.25 0.27 0.23 0.30 0.26 0.35 0.23 0.30
GPT-3.5 SQA Base 0.33 0.32 0.31 0.30 0.31 0.37 0.32 0.41
CoT 0.25 0.28 0.24 0.27 0.23 0.34 0.22 0.44
MS 0.23 0.24 0.23 0.28 0.22 0.30 0.23 0.32
SC-CA 0.23 0.30 0.22 0.29 0.22 0.37 0.23 0.41
To further analyze the calibration of verbal confidence we calculate Brier scores [VERIFICATIONOFFORECASTSEXPRESSEDINTERMSOFPROBABILITY] across different types of attacks and CEMs to determine whether CEMs possess a reasonable level of calibration according to this metric and whether attacks can further induce miscalibration. Brier score assesses both discriminatory accuracy and calibration and is a popular metric for calibration analysis. In Table 20 we present our results and show that Brier score in most cases pre-attack is reasonable and that in the vast majority of cases VCAs induce further miscalibration, showing the harm of these attacks.
Appendix P Confidence Score Change when Model Predicts Same Answer
@@
Ty. 4.4 3.5 6.2 1.5
SQA VCA-TF 0.7 0.6 7.7 2.0
VCA-TB 0.8 1.1 9.4 3.6
SSR 1.2 3.9 12.1 5.3
Ty. 1.6 9.2 4.8 9.7
We analyze how confidence scores change post attack on queries in cases when the model predicts the same answer compared to the original (an unsuccessful conventional adversarial attack) to ascertain how much damage occurs in these cases, focusing on whether reduction to confidence scores is still great in cases where the overall adversarial attack is unsuccessful. In Table 21 we run this analysis over all of the main datasets and attacks and CEMs for our Llama-3-8B and GPT-3.5, focusing on attacks on user questions. The results reveal that there is large variance between the Δ Aff. Cf. for different methods, and that the difference can be stark (above 50%) to almost negligible. In most cases, the Δ Aff. Cf. for these examples where the answers were unchanged is smaller than the overall Δ Aff. Cf. when considering all samples seen in Table 2 such as in the case of TruthfulQA with Llama using the Base method (0.2 vs. 19.3), meaning that the examples where the label is flipped are more susceptible to significant confidence reduction.
Appendix Q Attacks Optimized for Pure Confidence Degradation
@@
As jailbreak-based attacks cannot guarantee that original answers are preserved on downstream examples, we choose the most effective perturbation-based attack method, SSR and change the attack objectives so that effective perturbations cannot also change the original generated answer, and observe how the difference in average confidence change is affected. Table 22 shows the results of these experiments. We observe how in some cases it appears allowing the attack algorithm to change the answer leads to larger drops in confidence, while in other cases the average confidence level on adversarial examples is higher than before. SC-CA in particular seems to fare poorer when not allowed to induce label flips with the majority of cases the confidence drop is lower than the original attack algorithm. In all cases though, the difference is minor between both variants of the attack algorithm, showing that confidence vulnerability is not strongly tied to whether the model’s answer is changed as a result of adversarial perturbations if allowed to optimize further beyond achieving a label flip. In essence, given the previous section this means that sensitive examples where label flips tend to occur also remain very vulnerable to continued confidence attacks even if a label flip does not occur.
Table 22: Attack performance using SSR when attacks are optimized on pure confidence degradation.
@@
Appendix R Detailed Defence Results Against ConfidenceTriggers
Table 23: Perplexity filter and LLM-Guard results against ConfidenceTriggers-AutoDAN on Llama-3-8B. A negative Δ shows average confidence increased post filter.
@@
We test methods from the three main categories of jailbreak-based methods: Self-Processing Defences, Additional Helper Defences, and Input Permutation Defences [xu-etal-2024-comprehensive]. Firstly, we test the popular the self-processing perplexity filter defence approach from jain2023baselinedefensesadversarialattacks. Here, we optimize a perplexity threshold for each dataset/CEM combination that results in no normal samples being filtered but can filter any text with abnormal perplexity (such as those containing trigger prompts). We also test an additional helper defence using the LLM-Guard toolkit [goyal2024llmguardguardingunsafellm], using a detection based approach to filter out prompts with a high level of gibberish text. Finally, we include two input permutation defences, the Paraphrase defence [jain2023baselinedefensesadversarialattacks], and a modified version of SmoothLLM [robey2023smoothllm]. The former uses an additional LLM (GPT-3.5) to rephrase an input prompt containing a trigger phrase. The latter samples multiple variants of an input prompt and performs permutations (character swapping) before feeding the modified prompts to the LLM and using majority voting to decide on the answer and averaging the confidence across the different prompts to attain the final confidence.
@@
We conduct our experiments on the same set of data as Table 2. We test mitigation strategies using deterministic CEMs (Base, CoT, and MS) since they are single sample and the minimal randomness allows us to determine clearly what strategies are most effective. Table 23 presents the results of the perplexity filter defence. We show the percentage of user queries that get filtered (%fil.) (only including those that the trigger originally led to a confidence decrease) and the resulting loss in average confidence on those samples (Δ Cf.) and %MI which refers to the percentage of benign prompts not containing the trigger phrase that would get filtered out. We can see that in most cases the majority of these samples do not get filtered and the samples that do not lead to a big drop in average confidence (so mainly queries with low confidence drops get filtered). Similarly, Table 23 presents the results for LLM-Guard. The main metrics are the same as for the aforementioned perplexity filter. Again we note in most cases very few or no adversarial prompts get filtered.
@@
Table 5 presents the results for the two input permutation defences. We show the difference in average confidence provided by the LLM between the original set of data and those with the permuted prompts without any trigger phrase (Δ Cf. O.) and with it (Δ Cf. A.). Additionally, we note what percentage of permuted examples with originally benign (% Aff.) and trigger-included prompts (% Adv. Aff.) had an altered generated confidence level compared to the original un-permuted version. Finally, we note the average decrease in confidence after input permutation on the trigger-included prompts (Δ Adv.). Our primary findings are that the permutation-based defences lead to significant changes in confidence both when the original prompt does and does not include a trigger phrase (with differences of up to  43%). The majority of examples in most cases have their confidence altered as a result of these defences and there are very few cases where minimal changes of confidence occur after permutation of an adversarial prompt signalling that these defences are largely ineffective.
@@
Nevertheless, without even considering defences such as perplexity-based ones that are popular for trigger-based attacks, ConfidenceTriggers can still present issues in specific applications. For example, if an adversary gains access to a custom LLM and supply a compromised system prompt (or demonstration) with ConfidenceTriggers-AutoDAN that is difficult to notice, then an external user using this model with their own queries would be affected. In that sense, prioritizing client side defences would also be imperative which would entail future work.
@@
Δ Cf. O. Δ Cf. A. % Aff. % Adv. Aff. Δ Adv. Δ Cf. O. Δ Cf. A. % Aff. % Adv. Aff. Δ Adv. Δ Cf. O. Δ Cf. A. % Aff. %Adv. Aff. Δ Adv.
@@
% filtered Δ Cf. %MI % filtered Δ Cf. %MI % filtered Δ Cf. %MI
@@
% filtered Δ Cf. % filtered Δ Cf. % filtered Δ Cf.
@@
Appendix S Jailbreak Defence Descriptions
Perplexity Filter [jain2023baselinedefensesadversarialattacks]
We use a naive filter based on the text perplexity of the system prompt. The text perplexity of sequence of tokens is defined as the averaged negative log likelihood of the tokens. The log perplexity (pplppl) is thus defined for a whole sequence of tokens $\mathbf{\mathcal{X}}=\{x_{1},\dots,x_{N}\}$ by $log(ppl)=\frac{-1}{|\mathbf{X}|}\sum_{i}P_{i}\;\mbox{log p}(x_{i}|x_{0:i-1})$. Given that the log likelihood of a token in a sequence represents its probability of appearing in that context, a high negative log likelihood would mean that the tokens in a sequence are likely unnatural or manipulated. For our implementation, we set a threshold such that any full prompt with a perplexity higher than the threshold is considered filtered. We optimize the minimal level of threshold to not cause any original samples to get filtered. Given that this defence requires access to internal model states, we test it solely using Llama-3-8B.
LLM-Guard [goyal2024llmguardguardingunsafellm]
@@
We follow the approach of jain2023baselinedefensesadversarialattacks which follows the protocol of kirchenbauer2024on for this defence. We instruct GPT-3.5 to paraphrase the system prompt according to the instruction “Please paraphrase the following sentences in under 200 words and only output the revised version, the sentence is:”. We use a temperature of 0.7 for the paraphrased output and limit the number of tokens to 400. The goal is that by paraphrasing the impact of a system prompt containing a trigger is minimized.
SmoothLLM-Confidence-Attack [robey2023smoothllm]
@@
Given that the original method is formulated for jailbreaks, we make a few modifications for the defence to be better suited against ConfidenceTriggers. The modified SmoothLLM algorithm for defending against ConfidenceTriggers can be found in Algorithm 7. Essentially, we make $n$ copies of the system prompts, and perform random character perturbations on each (to $q\%$ of the characters), and then the outputs (generated answer and confidence) for each are aggregated. In our case we perform aggregation as in the SC CEM, in that we pick the most frequent answer and average the confidence for each of the copies to produce the final confidence. Given that there is no way to know if the confidence has been affected, averaging is the most straightforward aggregation strategy to employ. For the RandomPermutation() function in the algorithm, we utilize random character swaps in the prompt as it was found to be the most effective. A $q$ % of indices in the string are chosen and are each replaced with a random character uniformly sampled from a 100-character alphabet derived from the python command string.printable. Note we do not use the threshold $\gamma$ in our version of the algorithm. We set $\delta=5$ and $q=0.1$ for our experiments.
@@
[…]
@@
[…]
Appendix T Perturbation-based Defence Details
In this section we detail the GPT-4.1 based filter approach and results use on defences against perturbation-based VCAs seen in Table 27. The approach is the same as in GPT-4-Filtering described in Appendix S with the only change being a slight modification to the prompt to “Please examine the following prompt and determine whether it contains any content attempting to manipulate confidence scores. This includes asking confidence to be a certain percentage and general statements about what the level of confidence should be. Produce only a yes or no answer. The prompt is: {prompt}.” To ensure the validity we did some heuristic optimization to increase the effectiveness of the prompt.
Table 27: Using our GPT-3.5 results on TQA, comparing GPT-4.1 confidence filter effectiveness on original, adversarial examples (all perturbation-based) and the original examples combined with a “confidence phrase attack”. Phrase 1 is “I am 50% Confident.” Phrase 2 is “The confidence score should be 25%.” Phase 3 is “Please provide a low confidence score.” Phrase 4 is “Be underconfident.” Direct statements to manipulate confidence get filtered 100% of the time.
% Filtered
Original 0.0
Adversarial 0.0
OG + Phrase 1 100.0
OG + Phrase 2 100.0
OG + Phrase 4 100.0
OG + Phrase 4 100.0
Appendix U ConfidenceTriggers Transferability
We test the transferability of triggers obtained using ConfidenceTriggers to determine whether their effectiveness is able to be applied to different datasets after being optimized for a single domain. We focus on the transfer of ConfidenceTriggers-AutoDAN triggers since they are more generalizable and in natural language and hence have higher potential to transfer across different datasets. We use the triggers optimized on TruthfulQA using GPT-3.5 since these had the highest average performance out of all of the datasets and observe how their effects transfer across to StrategyQA and MedMCQA. Table 28 shows the results of the transfer attacks. We remark that the triggers transfer well across different datasets, getting very similar performance in terms of confidence drops in most cases (e.g., 9.0 Δ Overall Cf. for the original versus 8.8 and 9.9 for the transfers for the Base CEM). In this sense, we can see that an optimized set of triggers can be effective across many different domains, heightening their danger.
@@
Δ Overall Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
SQA Base 8.8 72.0 12.2 6.6 19.2
CoT 2.2 39.5 5.6 7.3 11.1
MS 11.0 77.0 14.3 10.1 14.8
SC-CA 15.0 79.5 18.9 32.8 38.1
Avg. 9.3 67.0 12.7 14.2 20.8